OpenStudy (anonymous):

Use the computational properties of logarithms to write the expression ln (x – 1)(x + 6) in another form.

OpenStudy (e.mccormick):

Know the rules of logs for things that are added and subtracted?

OpenStudy (anonymous):

No

OpenStudy (e.mccormick):

\[log(m\cdot n) = log(m)+log(n)\]That one applies here, if that log is around both of the terms.

OpenStudy (e.mccormick):

The other two are: \[log\left(\frac{m}{n}\right) = log(m)-log(n) \\ log(m^n)=n\cdot log(m)\]

OpenStudy (anonymous):

log(x-1)+log(x+6)?

OpenStudy (e.mccormick):

So if your is \(ln[(x-1)(x+6)]\) then yah, \(ln(x-1)+ln(x+6)\)

OpenStudy (anonymous):

and that's it?

OpenStudy (e.mccormick):

Not much you can do to what is inside the braces unless you could bring things to a power and then have the stuff simplify. Then your right had side would be e to the something and it should be solvable... but other than that, there is nothing else that could be done.

OpenStudy (e.mccormick):

This is the type of case where you could do more. \[ln x = 5\implies e^{ln x}=e^5 \implies x=e^5\] If they do not give you a right hand side, there is nothing else to be done.